IR Demo Project Weekly Report for May 10-14, 1999
Highlights for this week include successfully bringing the gun up to an adequate operating voltage (320 kV) with a cathode of improved quantum efficiency. Unfortunately the drive laser (after more than 2 years of highly reliable operation) degraded over last weekend to the point that it could not generate nominal electron beam; steps are being taken to repair it.
The FEL team fleshed out and priced the upgrade paths to 10-20 kW respective to the cases of incremental funding and full lump-sum funding.
FEL Installation/Maintenance Activities
The Happek bunch-length interferometer located in the injection line is being relocated to the entrance of the first recirculation arc to support coherent synchrotron radiation studies. The vacuum window at that location was replaced with one that is transmissive to infrared light, and the camera was installed there and aligned. The interferometer itself is being checked out and will soon be mounted there. Recall that we have never been able to obtain data with the "injector Happek". We believe that to be due to a too-long bunch length; we should find out as a spinoff of operating it in its new location.
FEL Commissioning Activities
Over the weekend we heat-cleaned the cathode, and on Monday we ran voltage up to 340 kV with the conditioning resistor in place. There were a few small discharges at 340 kV, but without any vacuum activity. At 335 kV field emission was completely quiet. Consequently we decided to run at 320 kV to ensure adequate overhead for operations and future recesiations. A cathode scan revealed about double the quantum efficiency over the previous two cathode wafers, so indications are that the gun performance would be fine, although a modest amount of extra time would be required to set up the injector given the gun's reduced voltage.
Unfortunately, over last weekend the drive laser power dropped by a factor of three so that it would not generate useful electron beam. Part of the weekend and the first half of the week were spent diagnosing the problem. The process led us to suspect failure of the rod, though we are not certain of the diagnosis. We contacted the vendor (Coherent in California) to request on-site technical support, and a representative will be here Monday with three new rods. By the time the drive laser is fixed and tuned, it will probably be Tuesday or Wednesday of next week before we can begin to run beam. That will leave about 8-9 days of run time before the planned shutdown of CEBAF's helium refrigerator, during which we should, at minimum, be able to reach the full kilowatt (or identify a "new" limitation) and take additional emittance data in support of CSR studies.